OpenStudy (caerus):

What value of capacitance will have a capacitive reactance of 80 ohms at 796 cycles?

OpenStudy (caerus):

@DanJS

OpenStudy (caerus):

@Luigi0210

OpenStudy (caerus):

@ganeshie8

OpenStudy (caerus):

@sammixboo

OpenStudy (osprey):

What value of capacitance will have a capacitive reactance of 80 ohms at 796 cycles? If "cycles" is the same thing as frequency, f, "Hertz", Hz ... reactance = 1/2 pi f C 2 pi f C = 1/reactance C=1/2 pi f C = 1/(2*3.142*796) 1.9994e-4 F -> 200 microfarads

OpenStudy (caerus):

theres an error , he didnt devide the reactance right?

OpenStudy (3mar):

I really don't know It is not from honesty to answer a question I do not know.

OpenStudy (caerus):

reactance=1/2 pi f C, therefore by solving for C... C=1/2 pi f(reactance) .... right? he forgot to plug the reactance which is 80 ohms... watch to his solution.

OpenStudy (3mar):

Right! So can you solve for C now?

OpenStudy (caerus):

yip

OpenStudy (3mar):

How?

OpenStudy (osprey):

apologies ... Just looking again ... C = 1/(2*3.142*796) C = 1/(2*3.142*796*80) apologies again

OpenStudy (3mar):

No problem, @osprey. It happens to all of us!

OpenStudy (osprey):

@3mar i usually put a disclaimer up and use words like "guess", but I've recently been warned for "spamming" or something, so now I'm pretty wary. Would be "interesting" to see what would happen if a circuit were constructed on the wrong calculation. Either it wouldn't work as intended - designer v annoyed with post, or it would explode, - designer v v annoyed with post ...